OpenStudy (anonymous):

A delivery truck is transporting boxes of two sizes: large and small. The large boxes weigh 50 pounds each, and the small boxes weigh 20 pounds each. There are 120 boxes in all. If the truck is carrying a total of 4050 pounds in boxes, how many of each type of box is it carrying?

OpenStudy (anonymous):

let x = number of large boxes and 120-x = number of small boxes 50x+20(120-x) = 4050 50x+2400-20x = 4050 30x = 1650 x=55 large boxes 120-x = 65 small boxes

OpenStudy (anonymous):

Suppose that there are two types of tickets to a show: advance and same-day. The combined cost of one advance ticket and one same-day ticket is 60$ For one performance, 25 advance tickets and 35 same-day tickets were sold. The total amount paid for the tickets was 1700$ . What was the price of each kind of ticket?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

You can set up a system of equations. Let x = the cost of an advance ticket and y = the cost of a same-day ticket. x + y = 60 25x+35y=1700 multiply the first equation by -25 -25x-25y=-1500 25x + 35y = 1700 Add the two together: 10y = 200 y = 20 x+20=60 x= 40 advanced tickets cost $40 and same day cost $20